OpenStudy (anonymous):

What is the interquartile range (IQR) for the data set? 13, 14, 17, 18, 23, 27, 28, 31, 34

OpenStudy (anonymous):

23

OpenStudy (anonymous):

Median = 23 1st quartile = (14+17)/2 = 15.5 3rd quartile = (28+31)/2 = 29.5 IQR = q3 - q1 = 29.5 - 15.5 = 14 therefore, the interquartile range is 14

Directrix (directrix):

@devanshgaur Code of conduct: OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

@Directrix ... OK I will be posting the solution.. I was justing asking katrina if I am right or not thats why I waited too long... I just wanted to consult.. Only then I will be posting the solution...

Directrix (directrix):

Okay, no worries. @devanshgaur